I'm having trouble with this problem it is 8x squared -16x+6=0. We are supposed to factor and solve it. Please help me understand it I have some of it worked out.
8x squared -16x+6=0 add: -16 multiply: 48
-1 times -48,-4 times -12
8x squared -4x-12x+6=0
I am stuck right here.

Found 2 solutions by solver91311, erika514:
Answer by solver91311(24713)



Ok, so far, so good.



Group into two binomials:



Go back and review that last step until you completely understand the reasons that the signs are the way they are, especially for the second binomial grouping.

Factor the GCF from each of the groups (4x from the first group, -6 from the second):



Notice that the binomials in parentheses are now identical. Factor them out.




From there it is just application of the Zero Product Rule.

John

Egw to Beta kai to Sigma
My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism


Answer by erika514(39)

Don't worry...we'll figure it out!;)
We need to factor our equation and once we do we should get 2(2x-1)(2x-3). We can get rid of the 2 out front by dividing by 2 and we still have (2x-1)(2x-3). Now we have two equations to solve for x and we will get two different solutions. For our first equation, we have 2x-1=0 so we add 1 to both sides to get x= 1/2. And for our other equation, we have 2x-3=0 so we add 3 to both sides and divide by 2 to get an answer of x=3/2. Our final solutions are x=1/2 and 3/2. Now we need to check to make sure these solutions work....I will do it for you this time, and they both work! In the future, the way you check your solutions is by plugging each value in for x and both sides of the equation should equal zero. Good luck, let me know if you have any more questions. :)